OpenStudy (cenaida):

Find the value for the variable expression h3 when h = 4. A. 7 B. 12 C. 36 D. 64

Mehek (mehek14):

h3 = h * 3 if h = 4 what is 4 * 3?

OpenStudy (cenaida):

12!

Mehek (mehek14):

Yes!

OpenStudy (cenaida):

thanks!!!!!

Mehek (mehek14):

your welcome

Nnesha (nnesha):

hmm i guess that is cube not h times 3

Nnesha (nnesha):

\[\huge\rm h^3\] like this

OpenStudy (cenaida):

yes it is

Nnesha (nnesha):

otherwise number suppose to be front of h

OpenStudy (cenaida):

like that

Nnesha (nnesha):

so replace h by 4

Mehek (mehek14):

ok so 4*4*4

OpenStudy (cenaida):

h^3

Nnesha (nnesha):

\[( 4)^3 = 4 \times 4 \times 4\]

OpenStudy (cenaida):

64?

Nnesha (nnesha):

yes that's right

OpenStudy (mikeymaximum):

I believe that its 12 as well

OpenStudy (cenaida):

k thanks guys :)

OpenStudy (mikeymaximum):

((thumbsup))

Mehek (mehek14):

its 64 @MikeyMaximum